the perimeter of a rectangle is 128. the length is 1 m more than twice the width. find the dimension, what is the length and the width
Answer by Maths68(1474)

the perimeter of a rectangle is 128. the length is 1 m more than twice the width. find the dimension, what is the length and the width


Length = l=2w+1 m
Width = w
Perimeter = p = 128 m


Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle =2 length +2 width
p = 2l +2w
Put the value of l and p in above equation
128 = 2(2w+1)+2w
128 = 4w+2+2w
128-2=6w
126=6w
Divide by 6 both sides of above equation
126/6=6w/6
21=w
w=21
Given
Length = l=2w+1=2(21)+1=42+1=43m

Dimension of rectangle
p=128
length=43
Width=21